# Asymptotic and compound e-values: multiple testing and empirical Bayes

### v2

@@ Abstract
We explicitly define the notions of (exact, approximate or asymptotic) compound p-values and e-values, which have been implicitly presented and extensively used in the recent multiple testing literature. While it is known that the e-BH procedure with compound e-values controls the FDR, we show the converse: every FDR controlling procedure can be recovered by instantiating the e-BH procedure with certain compound e-values. Since compound e-values are closed under averaging, this allows for combination and derandomization of FDR procedures. We then connect compound e-values to empirical Bayes. In particular, we use the fundamental theorem of compound decision theory to derive the log-optimal simple separable compound e-value for testing a set of point nulls against point alternatives: it is a ratio of mixture likelihoods. We extend universal inference to the compound setting. As one example, we construct approximate compound e-values for multiple t-tests, where the (nuisance) variances may be different across hypotheses. Finally, we provide connections to related notions in the literature stated in terms of p-values.
@@ 1 Introduction
The main purpose of this paper is to formalize the notions of asymptotic e-values, compound e-values and asymptotic compound e-values, show their versatility and usefulness for multiple testing, as well as provide certain optimal constructions and examples.
Contributions and Paper Outline.
In Section 2, we formally define approximate p-values and e-values, which are the building blocks of all contents in our paper, and provide several sets of conditions that lead to approximate p-values and e-values. In particular, Theorem 2.3 shows that calibration between exact p-values and exact e-values remains valid for approximate p-values and e-values, preserving their approximation parameters. This result will be useful in several constructions later in the paper. In Section 3, we introduce asymptotic p-values and e-values, connect them with usual limiting results in asymptotic statistics, and give a prototypical example of asymptotic e-values. In Section 4, we move to the setting of multiple testing, by defining compound e-values, studying their properties, and giving several examples. Their approximate and asymptotic versions are then discussed in Section 5, and a prototypical construction of approximate compound e-values for multiple t-tests is provided in Section 5.3. In Section 6, we describe the fundamental role of compound e-values in multiple testing: Theorem 6.2 (in Section 6.1) shows how to construct compound e-values from any FDR-controlling procedure, while Theorem 6.5 (in Section 6.2) demonstrates that every FDR procedure can be recovered by e-BH applied to compound e-values. We explain how this framework enables combination and derandomization of multiple testing procedures (Section 6.3) and that compound e-values provide a natural notion of weights for the weighted p-BH procedure (Section 6.4). In Section 7, we study sequence models and develop connections to compound decision theory, which motivates the nomenclature “compound e-values:” In Section 7.1, we provide a bird’s eye overview of compound decision theory, while in Section 7.2, we construct log-optimal simple separable compound e-values (Theorem 7.2 provides a formal statement). In Section 7.3, we introduce compound universal inference (CUI) and empirical Bayes universal inference (EBUI) for composite null hypotheses, proving in Proposition 7.4 that these constructions dominate universal inference e-values.
2 Approximate p-values and e-values
@@
Let $\mathbb{R}_{+}=[0,\infty)$. We say that a probability $\mathbb{P}$ is atomless if there exists a uniformly distributed random variable on $[0,1]$ under $\mathbb{P}$. In what follows, $x\wedge y=\min\{x,y\}$ and $x\vee y=\max\{x,y\}$.
@@
For two constants $\varepsilon\in\mathbb{R}_{+},$ and $\delta\in[0,1)$, $(\varepsilon,\delta)$-approximation is understood as those in Definition 2.1 with the corresponding constant functions.
Wang and Ramdas (2022, Section 6.2) considered $(\varepsilon,0)$-approximate e-variables (not using this terminology).
@@
Proof.
When $\mathbb{P}$ is atomless, for any random variable $X$, there exists an event $A$ with $\mathbb{P}(A)=1-\delta_{\mathbb{P}}$ such that $\{X<t\}\subseteq A\subseteq\{X\leq t\}$ for some $t\in\mathbb{R}$; see Lemma A.3 of Wang and Zitikis (2021). This fact will be used in the proof below. If $\delta_{\mathbb{P}}=1$ there is nothing to prove, so we assume $\delta_{\mathbb{P}}<1$ below.
(P2) $\Rightarrow$ (P1): It suffices to notice
for $\mathbb{P}\in\mathcal{P}$ and $t\in[0,1]$.
for all $t\in(0,1)$, as desired.
(E2) $\Rightarrow$ (E1): It suffices to notice
for $\mathbb{P}\in\mathcal{P}$ and $t\in\mathbb{R}_{+}$.
$\mathbb{E}^{\mathbb{P}}[E\mathds{1}_{A}]=\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A}]=\mathbb{E}^{\mathbb{P}}[E\wedge t]-\mathbb{E}^{\mathbb{P}}[(E\wedge t)\mathds{1}_{A^{c}}]\leq 1+\varepsilon_{\mathbb{P}}+t\delta_{\mathbb{P}}-\mathbb{E}^{\mathbb{P}}[t\mathds{1}_{A^{c}}]\leq 1+\varepsilon_{\mathbb{P}},$
as desired. ∎
@@
Given an e-variable $E$, we can calibrate it into a p-variable via $P=(1/E)\land 1$, which is the only admissible choice of an e-to-p calibrator. Conversely, given a p-variable $P$, we can calibrate into an e-variable via a (p-to-e) calibrator in the sense of Vovk and Wang (2021). A (p-to-e) calibrator is a decreasing function $h:[0,\infty)\to[0,\infty]$ satisfying $\int_{0}^{1}h(x)\mathrm{d}x\leq 1$ and $h(x)=0$ for $x\in(1,\infty)$. Given such $h$ and a p-variable $P$, then $E=f(P)$ is an e-variable. The next theorem demonstrates that we can calibrate approximate p-variables and approximate e-variables in exactly the same way.
@@
Proof..
$\leq(1+\varepsilon_{\mathbb{P}})\int_{0}^{1/b}f(s)\mathrm{d}s+tF(0)$
$\mathbb{P}(P\leq t)$ $=\mathbb{P}(E\geq 1/t)$
$=\mathbb{P}(E\wedge(1/t)\geq 1/t)$
as desired. ∎
Remark 2.4.
Following a similar proof, the statements in Theorem 2.3 hold true also for if we use the alternative formulations (P2) and (E2) in Proposition  2.2 for both approximate p-variables and approximate e-variables.
We end this section with the following comment. The roles of $(\varepsilon,\delta)$ are asymmetric and as mentioned earlier, the role of $\delta$ is more lenient. The following result shows that an $(\varepsilon,\delta)$-approximate p-variable (or e-variable) with $\varepsilon>0$ is always also an $(0,\delta^{\prime})$-approximate p-variable (or e-variable) where $\delta^{\prime}$ can be chosen strictly smaller than $\varepsilon+\delta$.
@@
Proof.
Meanwhile the inequality $\mathbb{P}(P\leq t)\leq t+\delta^{\prime}_{\mathbb{P}}$ is clearly true for $t\geq(1-\delta_{\mathbb{P}})/(1+\varepsilon_{\mathbb{P}})$ (since the right hand side is $\geq 1$). Thus $P$ is an $(0,\delta)^{\prime}$-approximate p-variable for $\mathcal{P}$.
Proposition 2.5 also suggests that controlling $\delta$ is a stronger requirement, leading to two different notions of asymptotic e-values and p-values in Section 3.
3 Asymptotic p-values and e-values
3.1 Definitions
@@
We would also say “asymptotic e-values” (or p-values) in non-mathematical statements, and they can be interpreted as asymptotic e-variables (or p-variables) or their realizations. We sometimes simply say that a sequence is asymptotic for $\mathcal{P}$, and it should be clear from context whether we meant a sequence of asymptotic e-variables ($[0,\infty]$-valued) or that of asymptotic p-variables ($\mathbb{R}_{+}$-valued).
Clearly, if $\mathcal{P}$ is finite, then asymptotic e-variables or p-variables are also uniformly asymptotic.
3.2 Conditions for asymptotic p-values and e-values
We view the requirement of being strongly asymptotic as particularly intuitive for e-variables due the following result.
@@ 3.2 Conditions for asymptotic p-values and e-values
Proof.
where the first inequality holds for large enough $n$. For the other direction, note that any nonnegative extended random variable $E^{(n)}$ is an $(\varepsilon_{n},0)$-approximate e-variable for $\mathcal{P}$ with the choice $\varepsilon_{n}(\mathbb{P})=\max\{0,\,\mathbb{E}^{\mathbb{P}}[E^{(n)}]-1\}$ as long as $\mathbb{E}^{\mathbb{P}}[E^{(n)}]$ is finite. If $\limsup_{n\to\infty}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1$, this implies that $\mathbb{E}^{\mathbb{P}}[E^{(n)}]<\infty$ for $n$ large enough and the above $\varepsilon_{n}(\mathbb{P})$ satisfies $\lim_{n\to\infty}\varepsilon_{n}(\mathbb{P})=0$, that is, $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of strongly asymptotic e-variables. ∎
@@
Proof.
Fix $\mathbb{P}\in\mathcal{P}$. Note that for $t\in\mathbb{R}_{+}$,
By uniform integrability, for any $\eta>0$ there exists $T>0$ such that
Thus,
and hence,
$\limsup_{n\to\infty}\mathbb{E}^{\mathbb{P}}[E^{(n)}]\leq 1+\eta.$
Since $\eta>0$ is arbitrary, we conclude by Proposition 3.2. ∎
@@
Proof.
By definition, if $\varepsilon:=f_{\delta}(E)\vee 0$ is finite, then $E$ is an $(\varepsilon,\delta)$-approximate e-variable for ${}\{\mathbb{P}\}$. Let $E$ be the limit of the nonnegative sequence $(E^{(n)})_{n\in\mathbb{N}}$ in distribution. Note that $f_{\delta}(E)\leq\mathbb{E}^{\mathbb{P}}[E]-1\leq 0$ for all $\delta\in[0,1]$. By the first-order condition, the supremum in $f_{\delta}(E)$ is attained for $t$ being a left or right $(1-\delta)$-quantile of $E$, and from there we can check that the function $f_{\delta}$ is continuous in convergence in distribution. This gives $\limsup_{n\to\infty}f_{\delta}(E^{(n)})\to f_{\delta}(E)\leq 0$. This implies, in particular, that for any $\delta>0$, there exists $N\in\mathbb{N}$ such that for $n>N$, $E^{(n)}$ is an $(\delta,\delta)$-approximate e-variable. Since $\mathbb{P}\in\mathcal{P}$ is arbitrary, this shows that $(E^{(n)})_{n\in\mathbb{N}}$ is a sequence of asymptotic e-variables for $\mathcal{P}$. ∎
@@
Gasparin et al. (2025) referred to asymptotic p-variables as those that converge to a p-variable in distribution. As Proposition 3.4 illustrates, this is a sufficient condition for our definition of asymptotic p-values. Our definition is more general. For instance, any sequence that is larger than a p-variable is an sequence of asymptotic p-variables, and it does not need to converge.
@@
Proof.
For (ii) we may argue as follows. If $h$ would have been continuous, then the result would follow from the continuous mapping theorem and Proposition 3.4. For upper semi-continuity, note that $\limsup_{n\to\infty}\mathbb{P}(E^{(n)}\geq x)\leq\mathbb{P}(h(Z)\geq x)$ for all $x\in\mathbb{R}$ by convergence in distribution. This gives $\limsup_{n\to\infty}f_{\delta}(E^{(n)})\leq f_{\delta}(h(Z))\leq 0$ using the notation and argument in the proof of Proposition 3.4, with some details omitted. The last statement follows from Proposition 3.3. When $h$ is increasing or decreasing, replacing it with its upper semi-continuous version, which does not change $\mathbb{E}[h(Z)]$, yields the desired statements. ∎
@@
The most common cases in practice include the following textbook examples (Van Der Vaart, 1998; Lehmann and Romano, 2005):
@@
3.3 A prototypical example of an asymptotic e-value
We next provide a prototypical example of asymptotic e-variables along the lines of the central limit theorem.
Example 3.9.
Suppose we observe the dataset $X^{(n)}=(X^{1},\dots,X^{n})$ where $X^{i}\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}$. We assume that their variance, $\mathrm{var}^{\mathbb{P}}(X^{i})$, is positive and finite, but is unknown. We seek to test the hypothesis
$\mathcal{P}=\{\mathbb{P}\in\mathcal{M}_{1}\,:\,\mathbb{E}^{\mathbb{P}}[X^{i}]=0,\,\mathrm{var}^{\mathbb{P}}({X^{i}})\in(0,\infty)\},$ (2)
where $\mathcal{M}_{1}$ denotes the set of all probability measures on $\mathbb{R}$. Define
$\bar{X}^{(n)}=\frac{1}{n}\sum_{i=1}^{n}X^{i},\;\;\;\,S^{(n)}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(X^{i})^{2}},\;\;\;\,\hat{\sigma}^{(n)}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}\left(X^{i}-\bar{X}^{(n)}\right)^{2}},$ (3)
and for fixed $\lambda\in\mathbb{R}$, let
and
The next proposition justifies that (4) defines a sequence of strongly asymptotic e-variables.
Theorem 3.10.
The sequences of random variables $E^{(n)}$ in (4) and $\tilde{E}^{(n)}$ in (5) are strongly asymptotic e-variables for $\mathcal{P}$ in (2). If $\hat{\sigma}^{(n)}$ is used in place of $S^{(n)}$ in (4) and (5), then $E^{(n)}$ and $\tilde{E}^{(n)}$ are asymptotic e-variables for $\mathcal{P}$. In either case, for $\lambda>0$, $E^{(n)}$ (resp. $\tilde{E}^{(n)}$) grows to infinity with probability $1$ under the alternative that $X^{1},\dots,X^{n}$ are drawn iid from $\mathbb{Q}$ with $\mathbb{E}^{\mathbb{Q}}[X^{i}]>0$ (resp. $\mathbb{E}^{\mathbb{Q}}[X^{i}]\neq 0$) and $\mathrm{var}^{\mathbb{Q}}(X^{i})<\infty$.
Proof..
for $Z\sim\mathrm{N}(0,1)$. By Proposition 3.8 it follows that $E^{(n)}$ are asymptotic e-variables. To upgrade this result to strongly asymptotic, we can apply Theorem 2.5 of Giné et al. (1997) to show that the sequence $(E^{(n)})_{n\in\mathbb{N}}$ is uniformly integrable and then conclude the proof by Proposition 3.3.
and
The above imply that
is a strongly asymptotic e-variable.
One central purpose of defining asymptotic e-variables is to handle cases where one does not know how to construct a nonasymptotic e-variable. Indeed, for the above $\mathcal{P}$, we do not know of nonasymptotic e-variables (other than constants). Had a bound on the variance been known, such nonasymptotic e-variables do exist, as shown next.
Example 3.11 (Testing the mean with bounded variance (Wang and Ramdas, 2023)).
Given some $\sigma>0$, consider the null hypothesis
$\mathcal{P}_{\sigma}=\left\{\mathbb{P}\in\mathcal{M}_{1}:\mathbb{E}^{\mathbb{P}}[X]=0,\;\mathrm{var}^{\mathbb{P}}({X})\leq\sigma^{2}\right\}.$
Then $X^{2}/\sigma^{2}$ is an e-variable for $\mathcal{P}_{\sigma}$, and for any $\lambda\in\mathbb{R}$ the following are e-variables for $\mathcal{P}_{\sigma}$:
$\exp\left(\lambda X-\frac{\lambda^{2}}{6}(X^{2}+2\sigma^{2})\right)$
and
$\exp\left(\phi(\lambda X)-\frac{\lambda^{2}}{2}\sigma^{2}\right),$
where $\phi(x)=\log(1+x+x^{2}/2)$ for $x\geq 0$ and $\phi(x)=-\log(1-x+x^{2}/2)$ for $x<0$. The different choices of e-variables have large e-power against different alternatives.
We leave the proof of the facts in Example 3.11 for the reader.
4 Compound p-values and e-values in multiple testing
We now extend the setting of Sections 2 and 3 to the multiple hypothesis testing setting. As before, we write $X$ to denote the entire dataset available to the researcher which is generated by some unknown probability distribution $\mathbb{P}^{*}$. We consider testing $K$ hypotheses determined by sets of probability measures $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$. For $k\in\mathcal{K}=\{1,\ldots,K\}$, the $k$-th hypothesis is defined as $H_{k}:\mathbb{P}^{*}\in\mathcal{P}_{k}$. We write
and hence $\mathcal{N}(\mathbb{P}^{*})$ denotes the indices of true null hypotheses.
We formally define the notion of compound p-variables and e-variables as fundamental statistical concepts. In what follows, $\mathcal{P}$ is a set of distributions that represents some assumptions or restrictions on the distributions in the testing problem.
Definition 4.1 (Compound p-variables and e-variables).
  1. (i)
  2. (ii)
$\mathcal{P}=\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k}.$
We can convert compound e-variables into compound p-variables and vice versa. If $E_{1},\dots,E_{K}$ are compound e-variables, then $(1/E_{1})\land 1,\dots,(1/E_{K})\land 1$ are compound p-variables, as can be seen by Markov’s inequality. Compound p-variables can be transformed into compound e-variables in the following result.
Theorem 4.2.
Proof.
$\tilde{P}:=P_{L}\mathds{1}_{\{U\leq K_{0}/K\}}+\mathds{1}_{\{U>K_{0}/K\}}$
is a bona fide p-value because for $t\in(0,1)$,
Therefore,
where the last inequality holds since $h$ is a p-to-e calibrator. ∎
Example 4.3 (Weighted p-variables and e-variables).
  1. (i)
  2. (ii)
compared to the condition for the compound p-variables/e-variables
Example 4.4 (Combining p-variables and compound e-variables).
Suppose that $P_{k}$ is a p-variable for $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$. $P_{1},\dots,P_{K}$ can be arbitrarily dependent. Moreover, let $E_{1},\dotsc,E_{K}$ be compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. If $(E_{1},\dots,E_{K})$ is independent of $(P_{1},\dots,P_{K})$, then $P_{1}/E_{1},\dots,P_{K}/E_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. We revisit this example in Section 6.4.
Example 4.5 (Convex combinations of compound e-variables).
Example 4.6 (Combining evidence from follow-up studies).
5 Approximate and asymptotic compound p/e-values

### v3

@@ Abstract
We explicitly define the notions of (bona fide, approximate or asymptotic) compound p-values and e-values, which have been implicitly presented and extensively used in the recent multiple testing literature. While it is known that the e-BH procedure with compound e-values controls the FDR, we show the converse: every FDR controlling procedure can be recovered by instantiating the e-BH procedure with certain compound e-values. Since compound e-values are closed under averaging, this allows for combination and derandomization of FDR procedures. We then connect compound e-values to empirical Bayes. In particular, we use the fundamental theorem of compound decision theory to derive the log-optimal simple separable compound e-value for testing a set of point nulls against point alternatives: it is a ratio of mixture likelihoods. As one example, we construct asymptotic compound e-values for multiple t-tests, where the (nuisance) variances may be different across hypotheses. Our construction may be interpreted as a data-driven instantiation of the optimal discovery procedure (ODP), and our results provide the first type-I error guarantees for data-driven ODP.
Keywords: Benjamini-Hochberg, false discovery rate, compound decision theory, optimal discovery procedure, nonparametric maximum likelihood
@@ 1 Introduction
E-values provide an alternative inferential device to p-values. E-values encompass likelihood ratios and have provided some of the first non-trivial tests for composite null hypotheses, e.g., for testing if the data distribution is log-concave (Dunn et al., 2025); see Ramdas and Wang (2024) for a comprehensive introduction to e-values. In a multiple testing setting with hypotheses indexed by $\mathcal{K}=\{1,\ldots,K\}$ and null indices $\mathcal{N}\subseteq\mathcal{K}$, e-values $E_{1},\ldots,E_{K}$ can be used alongside the eBH (e-Benjamini-Hochberg) procedure of Wang and Ramdas (2022) to control the false discovery rate (FDR) of Benjamini and Hochberg (1995). The defining property of the e-values $E_{1},\ldots,E_{K}$ is that they are $[0,\infty]$-valued random variables satisfying $\mathbb{E}[E_{k}]\leq 1$ for all $k\in\mathcal{N}$. In this paper, we study $[0,\infty]$-valued random variables $E_{1},\ldots,E_{K}$ satisfying the compound constraint that
$\sum_{k\in\mathcal{N}}\mathbb{E}[E_{k}]\leq K$ (1)
holds exactly in finite samples, or satisfying suitable approximate or asymptotic versions of this constraint. We call $E_{1},\ldots,E_{K}$ (bona fide, approximate, or asymptotic) compound e-values. Compound e-values (without being given this name) have played an important role in the recent multiple testing literature. As one notable example, Ren and Barber (2024) solved the open problem of derandomizing the model-X knockoff filter (Candès et al., 2018) for variable selection with FDR control by reinterpreting the procedure in terms of compound e-values (see Section 4.3 for more details on derandomization).
@@
The central aim of this paper is to set the statistical foundations for compound e-values, to demonstrate their central role, versatility, and usefulness for multiple testing, as well as provide certain optimal constructions and examples. In Section 2 we explicitly define compound e-values, study properties, and provide motivating examples. Our main contributions are as follows:
  • Compound e-values are central to multiple testing. In Section 4, we describe the fundamental role of compound e-values in multiple testing: Theorem 4.2 (in Section 4.1) shows how to construct compound e-values from any FDR-controlling procedure, while Theorem 4.5 (in Section 4.2) demonstrates that every FDR procedure can be recovered by e-BH applied to compound e-values. We explain how this framework enables combination and derandomization of multiple testing procedures (Section 4.3), including a novel application to derandomization of the randomized e-BH procedures of Xu and Ramdas (2024), and that compound e-values provide a natural notion of weights for the ep-BH procedure of Ignatiadis et al. (2024) (Section 4.4).
  • Sequence models and compound decisions. In Section 5, we study sequence models and develop connections to compound decision theory of Robbins (1951) that motivates the nomenclature “compound e-values.” In Section 5.1, we provide a bird’s eye overview of compound decision theory, while in Section 5.2, we construct log-optimal simple separable compound e-values using the argument underlying the fundamental theorem of compound decision theory (Theorem 5.2 provides a formal statement). These optimal compound e-values are a ratio of mixture likelihoods and their form is identical to the statistics of the optimal discovery procedure (ODP) of Storey (2007). For this reason, we call them optimal discovery compound e-values.
[…]
  • Approximate and asymptotic compound e-values in multiple testing. While there are very few results on multiple testing with asymptotic and/or compound p-values, we show that the analogous notions with e-values are naturally compatible with guarantees for existing multiple testing procedures. Section 4 provides this unifying perspective. This is why we present Section 4 (on the central role of compound e-values in multiple testing) after Section 3 (which defines approximate and asymptotic compound e-values). In Section 7 we provide an alternative notion of approximate compound e-values that we call *approximate compound e-values. While these definitions coincide for single e-values, they differ in the compound case. The alternative definition is easier to verify but less compatible with existing multiple testing procedures.
  • Localized and compound universal inference. In Section 5.3, we extend universal inference (UI) of Wasserman et al. (2020) for testing composite null hypotheses and introduce localized UI (LUI) and compound UI (CUI) universal inference. LUI is a general construction leading to approximate e-values, while CUI leads to approximate compound e-values.
  • The simultaneous t-test sequence model. In Section 6 we demonstrate the usefulness of our definitions and constructions by instantiating them in the canonical setting of multiple t-tests where the nuisance parameters (variances) may be different across hypotheses. We first consider the normal setting (normal errors) with fixed sample-size for each hypothesis but growing number of hypotheses. Section 6.1.1 provides an explicit construction of CUI e-values, while Section 6.1.2 approximates the optimal discovery compound e-values using empirical Bayes, specifically, by pretending the variances are iid from a prior $G$ and estimating $G$ via the nonparametric maximum likelihood estimator (NPMLE) of Robbins (1950) and Kiefer and Wolfowitz (1956). Theorem 6.3 shows that this construction yields asymptotic compound e-values as $K\to\infty$. Our construction can be interpreted as a data-driven implementation of ODP, and to the best of our knowledge, Theorem 6.3 provides the first type-I error guarantee for data-driven ODP. The usefulness of these procedures is illustrated via a simulation study in Section 6.1.3 wherein we demonstrate substantial power-gains compared to alternatives such as UI and standard t-test e-values. In Section 6.2 we provide constructions that do not require normality: we provide a prototypical example of an asymptotic e-value (Section 6.2.1) that requires the sample size to grow to $\infty$, and then (Section 6.2.2) we provide an illustrative example of how to construct asymptotic compound e-values when either the per-hypothesis sample size or the number of hypotheses (or both) tends to $\infty$.
  • Connections to p-values. Throughout the paper, we establish connections between our definitions and results and their analogous counterparts for p-values. In Section 2 we define compound p-values, which have been previously called average significance controlling p-values by Armstrong (2022). Section 3 introduces approximate and asymptotic compound p-values, while Section 7 presents *approximate compound p-values; both of these concepts are new. A key finding is that the calibration between p-values and e-values established by Vovk and Wang (2021) extends to all variants introduced in this paper. We establish this correspondence across compound (Theorem 2.3), approximate (Theorem 3.3), asymptotic (Proposition 3.11), approximate compound (Proposition 3.14), and *approximate cases (Proposition 7.3). We also discuss results on multiple testing with compound p-values in Section 4.5, specifically generalizing a result of Armstrong (2022) on the Benjamini-Yekutieli procedure (BY) (Benjamini and Yekutieli, 2001) with compound p-values.
Remarks on notation and terminology: We let $\mathbb{R}_{+}=[0,\infty)$. We say that a probability measure $\mathbb{P}$ is atomless if there exists a uniformly distributed random variable on $[0,1]$ under $\mathbb{P}$. We also write $x\wedge y=\min\{x,y\}$, $x\vee y=\max\{x,y\}$, and $x_{+}=x\vee 0$.
2 Compound p-values and e-values in multiple testing
@@
We now extend the setting above to multiple hypothesis testing. As before, we write $X$ to denote the entire dataset available to the researcher which is generated by some unknown probability distribution $\mathbb{P}^{*}$. We consider testing $K$ hypotheses determined by sets of probability measures $\mathcal{P}_{1},\ldots,\mathcal{P}_{K}$. For $k\in\mathcal{K}=\{1,\ldots,K\}$, the $k$-th hypothesis is defined as $H_{k}:\mathbb{P}^{*}\in\mathcal{P}_{k}$. We write
and hence $\mathcal{N}(\mathbb{P}^{*})$ denotes the indices of true null hypotheses. In the multiple testing setting, we use the notation $\mathcal{P}$ for a set of distributions that represents some assumptions or restrictions on the distributions in the multiple testing problem.
We formally define compound p-variables and e-variables as fundamental statistical concepts.
Definition 2.1 (Compound p-variables and e-variables).
  1. (i)
  2. (ii)
$\mathcal{P}=\bigcup_{k\in\mathcal{K}}\mathcal{P}_{k}.$
Example 2.2 (Weighted p-variables and e-variables).
  1. (i)
  2. (ii)
compared to the condition for the compound p-variables/e-variables
Given an e-variable $E$, we can calibrate it into a p-variable via $P=(1/E)\land 1$, which is the only admissible choice of an e-to-p calibrator. Conversely, given a p-variable $P$, we can calibrate into an e-variable via a (p-to-e) calibrator in the sense of Vovk and Wang (2021). A (p-to-e) calibrator is a decreasing function $h:[0,\infty)\to[0,\infty]$ satisfying $\int_{0}^{1}h(x)\mathrm{d}x\leq 1$ and $h(x)=0$ for $x\in(1,\infty)$. Given such $h$ and a p-variable $P$, then $E=f(P)$ is an e-variable. We can analogously convert compound e-variables into compound p-variables and vice versa. If $E_{1},\dots,E_{K}$ are compound e-variables, then $(1/E_{1})\land 1,\dots,(1/E_{K})\land 1$ are compound p-variables, as can be seen by Markov’s inequality. Compound p-variables can be transformed into compound e-variables in the following result.
Theorem 2.3.
Example 2.4 (Combining p-variables and compound e-variables).
Suppose that $P_{k}$ is a p-variable for $\mathcal{P}_{k}$ for all $k\in\mathcal{K}$; $P_{1},\dots,P_{K}$ can be arbitrarily dependent. Moreover, let $E_{1},\dotsc,E_{K}$ be compound e-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. If $(E_{1},\dots,E_{K})$ is independent of $(P_{1},\dots,P_{K})$, then $P_{1}/E_{1},\dots,P_{K}/E_{K}$ are compound p-variables for $(\mathcal{P}_{1},\dots,\mathcal{P}_{K})$ under $\mathcal{P}$. We revisit this example in Section 4.4.
Example 2.5 (Convex combinations of compound e-variables).
Example 2.6 (Combining evidence from follow-up studies).
3 Approximate/asymptotic compound p-values and e-values
This section relaxes Definition 2.1 to accommodate approximate and asymptotic variants of compound p-variables and e-variables. We begin by introducing the underlying concepts of approximate p-variables and e-variables (Section 3.1) and asymptotic p-variables and e-variables (Section 3.2), then extend these to the compound setting in Section 3.3 (approximate compound) and Section 3.4 (asymptotic compound).
3.1 Approximate p-values and e-values
@@
For two constants $\varepsilon\in\mathbb{R}_{+}$ and $\delta\in[0,1]$, $(\varepsilon,\delta)$-approximation is understood as those in Definition 3.1 with the corresponding constant functions.
The notion of $(\varepsilon,0)$-approximate e-variables (not using this terminology) was considered by Wang and Ramdas (2022, Section 6.2).
@@
The next theorem demonstrates that we can calibrate approximate p-variables and approximate e-variables in the usual way. For example, if $h$ is any p-to-e calibrator and $P$ is an $(\varepsilon,\delta)$-approximate p-variable, then $h(P)$ is an ($\varepsilon,\delta)$-approximate e-variable.
@@
We end this section with the following comment. The roles of $\varepsilon$ and $\delta$ are asymmetric and, as mentioned earlier, the wiggle room created by $\delta$ is more lenient. The following result shows that an $(\varepsilon,\delta)$-approximate p-variable (or e-variable) with $\varepsilon>0$ is always also an $(0,\delta^{\prime})$-approximate p-variable (or e-variable) where $\delta^{\prime}$ can be chosen strictly smaller than $\varepsilon+\delta$.
@@
Proposition 3.4 also suggests that controlling $\delta$ is a stronger requirement, leading to two different notions of asymptotic e-values and p-values in Section 3.2.
3.2 Asymptotic p-values and e-values
@@
We would also say “asymptotic e-values” (or p-values) in non-mathematical statements, and they can be interpreted as asymptotic e-variables (or p-variables) or their realizations. We sometimes simply say that a sequence is asymptotic for $\mathcal{P}$, and it should be clear from context whether we meant a sequence of asymptotic e-variables ($[0,\infty]$-valued) or that of asymptotic p-variables ($\mathbb{R}_{+}$-valued).111 Liu and Singh (1997) use the terminology “limiting p-values” for asymptotic p-values and “limiting p-values in the strong sense” for uniformly asymptotic p-values. Clearly, if $\mathcal{P}$ is finite, then asymptotic e-variables or p-variables are also uniformly asymptotic.
We now study conditions and properties of asymptotic p-variables and e-variables. Our first result offers intuition regarding the formulation of strongly asymptotic e-variables.
@@
Gasparin et al. (2025) referred to asymptotic p-variables as those that converge to a p-variable in distribution. As Proposition 3.8 illustrates, this is a sufficient condition for our definition of asymptotic p-values. Our definition is more general. For instance, any sequence that is larger than a p-variable is a sequence of asymptotic p-variables, and it does not need to converge.
@@
The most common cases in practice include the following textbook examples (Van Der Vaart, 1998; Lehmann and Romano, 2005) (also see Section 6.2.1 for an explicit example):